DYNAMICS OF QUADRATIC POLYNOMIALS:

COMPLEX BOUNDS FOR REAL MAPS

Mikhail Lyubich Michael Yampolsky

Abstract. We extend Sullivan’s complex a priori bounds to real quadratic polynomials with essentially bounded combinatorics. Combined with the previous results of the first author, this yields complex bounds for all real quadratics. Local connectivity of the corresponding Julia sets follows.

Typeset by AMS-TEX